# Recent Advances and Trends in Research Paper Recommender Systems: A Comprehensive Survey

### v1

@@ 3.1.5 How do the RP-RSs generate the recommendations?
Inner product - Cross-entropy loss - [30, 68, 17, 40, 75]
Tuple-wise loss - [81]
WMS loss - [32]
- - [21, 74, 24, 69, 51, 25, 3, 66, 57]
BRP loss - [20]
BPR loss - [55, 18, 67]
Cutomized loss - [16]
WARP loss - [1]
MSE loss - [60]
Similarity Cosine similarity, Jaccard similarity - - [53]
Cosine similarity Hierarchical loss - [42]
Contrastive loss - [63]
- - [22, 43, 46, 11, 52, 38, 62, 49]
Cosine similarity, Citation similarity - - [58]
JSD , PageRank - - [23]
Cosine similarity, Clustering - - [72]
Cosine limilarity, Citation similarity - HoIN(CN) [39]
Meta-path, RW, PathSim - HIN [76]
JSD - - [13]
Term Frequency-based - - [44]
ML-based NN Cross-entropy loss - [78, 71, 65, 79]
Customized loss - [77, 70]
MSE loss - [64]
- - [54]
Memory network Cross-entropy loss - [4]
Auto-Encoder Customized loss - [45]
RNN, SVM - [15]
GCN, Outer product BPR loss - [73]
Graph-based Meta-path, RW BPR loss HIN [61]
COOT algorithm - CN [56]
Meta-path, RWR - HIN [29, 28]
RWR - HIN [26]
Other Co-citation, Bibliographic coupling - HoIN(CN) [31]
Personalized(NaÏve Bayes classifier/Softmax) - - [33]
Personalized score, PPR - - [80]
Bibliographic coupling, Similarity - - [19]
Hybrid(Relevance score + Inner product) - - [50]
@@ 3.2.1 ACL dataset
[77] ? ur Papers: 18,743 Authors: 14,455 Venues: 247
@@
[4] AAN Papers: 21,450 Authors: 17,335 Venues: 311
@@
[43] AAN Papers: 16,664 Authors: 13,934 Citations: 81,379
[20] AAN  [^7] Extracted from publicly available subdataset https://github.com/AHULiuYang/WHIN-CSL (Release date: 2013-02) Papers: 13,375 Authors: 3,936 Citations: 56,002
@@
[26] AAN Papers: 13,426 Authors: 10,927 Venues: 297
@@ 3.2.2 DBLP
[17] DBLP Papers: 1,201,716 Authors: 701,923
@@
[16] DBLP Papers: 14,376 Authors: 14,475
@@
[72] DBLP Papers: 21,044 Authors: 28,646
@@
[61] DBLP Papers: 2,126,267 Users: 3,765
@@
[78] DBLP ur Papers: 60,034 Authors: 196,957
@@
[69] DBLP-CN-? Papers: 10,291 User(Author): 17,616
@@
[18] DBLP-CN-? Papers: 38,714 Users(Authors): 26,941
@@
[68] DBLP-CN-? Papers: 9,566 Users(Authors): 16,863
@@
[68] DBLP-CN-? Papers: 28,930 Users(Authors): 7,126
@@
[23] DBLP-CN-V4 Large-scale dataset: Subset:
@@
[63] DBLP-CN-V10  [^10] https://github.com/opendata-ai/tr Papers: 21,000 Years: 1967-2017
[30] DBLP-CN-V11 Papers: 23,012 s-r: 1,037
@@
[55] DBLP-CN-V11 Papers: 17,414 Users(Authors): 3,866
@@
[77] DBLP-CN-V11 ur Papers: 3,212,633 Authors: 219,732
[70] DBLP-CN-V11 ur Papers: 716,549 Authors: 976,544
@@
[4] DBLP-CN-V12 Papers: 3,501,133 Citations: 25,022,314
@@
[20] DBLP-CN-V12 Papers:14,545 Authors:3,768
@@
[26] DBLP-CN-V12 Papers: 60,407 Authors: 77,172
@@
[55] DBLP-CN-V13 Papers: 51,686 Users(Authors): 12,068
@@
A significant number of works have used published versions of this dataset collection (see Table LABEL:tab:DBLPDataset) to conduct experiments without performing any data extraction or modifications. These include DBLP-CN-V12 [31, 64, 65] and DBLP-CN-V13 [56, 64, 57, 65]
@@ 3.2.3 AMiner
[77] ? ur Papers: 3,499,032 Authors: 325,341
@@
[16] ? Papers: 500,000 Keywords: 9,668
[80] ASN + Scopus  [^13] https://github.com/Pratyush1296/SPACE-R Papers: 12,611 Authors: 22,603
@@
[21] ASN + AFT [^14] https://academictree.org/ Papers: 89,013 Students: 10,929
@@ 3.2.4 CiteULike
Among the works analyzed, 26.98% utilize some CiteULike version, yet these represent only 8.3% of the total datasets used, showing a greater reusability of the CiteULike dataset compared to others. In fact, almost all works that use CiteULike rely on CiteULike-a [25, 73, 75, 3, 74, 18, 24, 32, 1, 80, 60] and CiteULike-t [25, 73, 75, 74, 18, 24, 32, 60] without applying any filtering or modification.
@@
[67, 66] ? ur Papers: 82,376 Users: 1,659
@@
[81] CiteULike-a Papers:11,845 Users: 4,880
@@
[15] CiteULike-a Papers: 1,000 Users: 10,327
@@
[28, 29] CiteULike-a 171717ut-kdd/CiteUlike-SPIN-dataset(github.com) Papers: 16,980 Users: 5,551
@@ 3.2.5 Other source datasets
In addition to the previously mentioned datasets, there exists another group of publicly available datasets that, while published, are less frequently utilized in the analyzed works. Of the works analyzed, 17.46% utilize some of these datasets, representing 13.89% of the total datasets used; however, only three of these works employ the datasets in their original, unmodified form [46, 49, 13].
@@
[46, 49] Sugiyama [^18] http://scholarbank.nus.edu.sg/handle/10635/146027[34, 59] Researchers: 50 Candidate Papers: 100,351
@@
[79] US patent [^21] https://developer.uspto.gov/data Papers: 182,260 Authors: 73,974
@@
[42] PWC [^22] https://github.com/malteos/aspect-document-embeddings[48] ur Papers: 88,940 Disciplines: Machine
@@
[42] SCICITE [^23] https://github.com/allenai/scicite ur Papers: 8,045 Disciplines:
@@
[69] Spotify [^24] https://www.kaggle.com/datasets/andrewmvd/spotify-playlists Papers: 11,703 Users: 1,720
[69] Yelp [^25] https://www.kaggle.com/datasets/yelp-dataset/yelp-dataset Papers: 19,479 Users: 10,918
[16] Metallurgical ? Papers: 5,702 Authors: 3,268
@@
[43] Arxiv HEP-TH [^26] https://snap.stanford.edu/data/cit-HepTh.html Papers: 18,066 Authors: 11,879
@@
[53] ? ur Disciplines: Computer Science
[58] ? Papers: 2,271 Disciplines: Carbon,
@@
[39] DBLP Papers: 700
[76] CNKI [^33] https://eval.cnki.net/index/ ur Disciplines:
@@
[76] CNKI ur Disciplines: Data Mining
@@
[80] WoS  [^34] https://github.com/Pratyush1296/SPACE-R Papers: 27,442 Authors: 40,408
@@
[33] PubMed  [^35] https://github.com/bioinfcollab/emati Disciplines:
@@
[11] Springer Nature
@@
[38] ArXiv Papers: 41,000 Users: 24,707
@@
[22] ArXiv Papers: 122,825 Disciplines: Computer science
[63] ArXiv  [^36] https://github.com/opendata-ai/tr. Papers: 21,000 Years: 1991-2016
[79] Scopus Papers: 1,304,907 Authors: 482,602
@@
[15] Scopus Size: 2000 Authors: 430
@@
[62] Semantic Scholar Papers: 805,063 Disciplines: Engineering,
@@
[52] Semantic Scholar +  [^37] https://opendata.iisys.de/opendata/Datasets/publication_similarity.zip Papers: 15,359 Publication Pairs:
@@
[50] Semantic Scholar +  [^38] https://github.com/IratxePinedoehu/arzigo Papers: 1795 Users: 736
@@ 3 Medicine
[71] ACM Papers: 3,025 Authors: 5,835
@@
[63] USPTO  [^39] https://github.com/opendata-ai/tr. Papers: 21,000 Years: 2002-2018
@@ 3.3.1 What are the relevance criteria used to evaluate the recommendations?
44.29% of the works use references as a relevance criterion. 38.71% of them consider the author as the user, so the papers referenced by authors in their publications are used as a basis for relevance [55, 68, 45, 79, 17, 16, 69, 77, 51, 21, 30, 20]. 38.71% of the works using references as relevance criteria rely on references found in the input paper [54, 65, 43, 39, 23, 58, 80, 4, 22, 26, 64, 63]. 12.9% of the works using references, even if the authors are not established as users, use datasets that only consider citation information for evaluating the system, so authors are considered the users, and their references are treated as interactions [61, 18, 15, 67]. 6.45% of the works use references from the same paper that is used to extract the input text [62, 56]. [70] not only considers positive examples the articles referenced by each author, but also the papers they have published.
@@
Out of the total works that use actual user interactions as a relevance criterion, 85% utilize the CiteULike dataset. The remaining works either use non-RPRS-specific datasets, such as Spotify or Yelp [69], or rely on artificially generated interactions [57, 33].
@@
12.86% of the works use some kind of label to identify relevant papers. These labels can be discipline classification labels that help match academic fields between users and recommended papers [40, 71] or between POI and recommended papers [38, 13]. Citation intent, which reflects the task or focus of the paper, is also used [42]. In addition, more explicit labels are sometimes applied to indicate relevance to particular authors [49, 49] or topics [19].
@@
Only 8.57% of the works use human judgment to evaluate the recommendations obtained. This relevance criterion can help assess the recommendations by better aligning them with actual needs, leading to more accurate evaluations. The choice of criteria depends on the level of detail required and the specific context in which the evaluation is conducted. Some systems use simple binary ratings, such as categorizing publications as either relevant or irrelevant [44], or correct or incorrect with reasoning [50]. Others implement more nuanced scales, assigning points based on similarity or relevance strength [52, 11, 53]. In addition to rating systems, some evaluations involve assessing the satisfaction of the users, highlighting the subjective nature of relevance judgment [31].
@@
Finally, alternative relevancy criteria in evaluating recommendations include assessing co-authorship potential by evaluating research compatibility [78], determining success based on the inclusion of papers by scientific research users [72], and matching recommendations with user input parameters such as research topics and academic affiliations [76].
@@
Paper references [54, 62, 55, 68, 45, 79, 18, 17, 65, 16, 70, 67, 56, 43, 69, 77, 39, 23, 51, 58, 15, 80, 21, 4, 22, 26, 30, 64, 20, 61, 63] 44.29%
Interactions [32, 1, 60, 28, 81, 18, 67, 29, 3, 33, 69, 74, 15, 80, 24, 57, 75, 66, 73, 25] 28.57%
Labeled Paper [49, 40, 71, 38, 13, 42, 46, 19, 63] 12.86%
Human judment [52, 11, 53, 50, 31, 44] 8.57%
Others [72, 78, 76] 4.29%

### v2

@@ 3.1.5 How do the RP-RSs generate the recommendations?
Inner product - Cross-entropy loss - [IJ:Scientometrics:Jiang:2023, IC:DASFAA:Wang:2022, IJ:AI:GAO:2023, CF:CSCW:LI:2022, IJ:SC:XI:2024]
Tuple-wise loss - [IC:KSEM:YU:2022]
WMS loss - [arXiv:KARIMI:2022]
- - [IJ:mathematics:GUO:2024, IJ:ESWA:XI:2023, IJ:JS:Huang:2023, IJ:ESWA:Wang:2023, IJ:IEEE:REN:2022, J:EIJ:Huang:2024, IJ:CDS:AHMEDI:2022, IJ:SMC:WANG:2022, IJ:KIS:STERGIOPOULOS:2024]
BRP loss - [IJ:TIIS:GUO:2024]
BPR loss - [IC:ASSP:Shen:2024, IC:MLBDBI:GUAN:2022, IJ:ASC:Wang:2024]
Cutomized loss - [IJ:AS:Chen:2023]
WARP loss - [ARXIV:Mohamed:2022]
MSE loss - [IC:ICITACEE:Switrayana:2022]
Similarity Cosine similarity, Jaccard similarity - - [IJ:IEEE:SARWAR:2021]
Cosine similarity Hierarchical loss - [IJ:IPM:Long:2024]
Contrastive loss - [IJ:Scientometrics:TAN:2023]
- - [IJ:Scientometrics:GUNDOGAN:2022, IJ:ESWA:MEI:2022, IJ:SECS:Hamisu:2024, IJ:ELECTRONICS:Blazevic:2023, IC:WIST:Roßner:2023, IJ:AMEC:KUS:2023, EC:ECIR:Takahashi:2022, AC:ACIIDS:PAN:2024]
Cosine similarity, Citation similarity - - [IJ:IEEE:STERLING:2022]
JSD , PageRank - - [IJ:IEEE:HADHIATMA:2023]
Cosine similarity, Clustering - - [IJ:JES:LI:2024]
Cosine limilarity, Citation similarity - HoIN(CN) [IJ:Heliyon:LI:2024]
Meta-path, RW, PathSim - HIN [IJ:AI:Xiao:2023]
JSD - - [IJ:EIT:Smail:2023]
Term Frequency-based - - [J:JIS:MUNGEN:2022]
ML-based NN Cross-entropy loss - [IJ:KBS:Xiao:2023, IC:SEAI:Wang:2024, IJ:AI:Thierry:2023, IC:ICDE:XIE:2022]
Customized loss - [IJ:ESWA:XIAO:2023, IJ:AS:Wang:2024]
MSE loss - [IJ:Scientometrics:Thierry:2023]
- - [ARXIV:Shen:2024]
Memory network Cross-entropy loss - [IJ:Scientometrics:ALI:2022]
Auto-Encoder Customized loss - [IC:EITCE:NIU:2023]
RNN, SVM - [IJ:IS:Chaudhuri:2022]
GCN, Outer product BPR loss - [IS:CSET:WU:2022]
Graph-based Meta-path, RW BPR loss HIN [$P$:Mahesh:2023]
COOT algorithm - CN [IJ:DSM:Sivasankari:2023]
Meta-path, RWR - HIN [IJ:AUT:Jafari:2023, IC:IKT:JAFARI:2021]
RWR - HIN [IJ:Scientometrics:Huang:2024]
Other Co-citation, Bibliographic coupling - HoIN(CN) [IJ:Scientometrics:Kanwal:2024]
Personalized(NaÏve Bayes classifier/Softmax) - - [IJ:DATABASE:Kart:2022]
Personalized score, PPR - - [IJ:JOI:Yadav:2022]
Bibliographic coupling, Similarity - - [IJ:SHTI:GUO:2022]
Hybrid(Relevance score + Inner product) - - [IJ:IS:PINEDO:2024]
@@ 3.2.1 ACL dataset
[IJ:ESWA:XIAO:2023] ? ur Papers: 18,743 Authors: 14,455 Venues: 247
@@
[IJ:Scientometrics:ALI:2022] AAN Papers: 21,450 Authors: 17,335 Venues: 311
@@
[IJ:ESWA:MEI:2022] AAN Papers: 16,664 Authors: 13,934 Citations: 81,379
[IJ:TIIS:GUO:2024] AAN  [^7] Extracted from publicly available subdataset https://github.com/AHULiuYang/WHIN-CSL (Release date: 2013-02) Papers: 13,375 Authors: 3,936 Citations: 56,002
@@
[IJ:Scientometrics:Huang:2024] AAN Papers: 13,426 Authors: 10,927 Venues: 297
@@ 3.2.2 DBLP
[IJ:AI:GAO:2023] DBLP Papers: 1,201,716 Authors: 701,923
@@
[IJ:AS:Chen:2023] DBLP Papers: 14,376 Authors: 14,475
@@
[IJ:JES:LI:2024] DBLP Papers: 21,044 Authors: 28,646
@@
[$P$:Mahesh:2023] DBLP Papers: 2,126,267 Users: 3,765
@@
[IJ:KBS:Xiao:2023] DBLP ur Papers: 60,034 Authors: 196,957
@@
[IJ:ESWA:Wang:2023] DBLP-CN-? Papers: 10,291 User(Author): 17,616
@@
[IC:MLBDBI:GUAN:2022] DBLP-CN-? Papers: 38,714 Users(Authors): 26,941
@@
[IC:DASFAA:Wang:2022] DBLP-CN-? Papers: 9,566 Users(Authors): 16,863
@@
[IC:DASFAA:Wang:2022] DBLP-CN-? Papers: 28,930 Users(Authors): 7,126
@@
[IJ:IEEE:HADHIATMA:2023] DBLP-CN-V4 Large-scale dataset: Subset:
@@
[IJ:Scientometrics:TAN:2023] DBLP-CN-V10  [^10] https://github.com/opendata-ai/tr Papers: 21,000 Years: 1967-2017
[IJ:Scientometrics:Jiang:2023] DBLP-CN-V11 Papers: 23,012 s-r: 1,037
@@
[IC:ASSP:Shen:2024] DBLP-CN-V11 Papers: 17,414 Users(Authors): 3,866
@@
[IJ:ESWA:XIAO:2023] DBLP-CN-V11 ur Papers: 3,212,633 Authors: 219,732
[IJ:AS:Wang:2024] DBLP-CN-V11 ur Papers: 716,549 Authors: 976,544
@@
[IJ:Scientometrics:ALI:2022] DBLP-CN-V12 Papers: 3,501,133 Citations: 25,022,314
@@
[IJ:TIIS:GUO:2024] DBLP-CN-V12 Papers:14,545 Authors:3,768
@@
[IJ:Scientometrics:Huang:2024] DBLP-CN-V12 Papers: 60,407 Authors: 77,172
@@
[IC:ASSP:Shen:2024] DBLP-CN-V13 Papers: 51,686 Users(Authors): 12,068
@@
A significant number of works have used published versions of this dataset collection (see Table LABEL:tab:DBLPDataset) to conduct experiments without performing any data extraction or modifications. These include DBLP-CN-V12 [IJ:Scientometrics:Kanwal:2024, IJ:Scientometrics:Thierry:2023, IJ:AI:Thierry:2023] and DBLP-CN-V13 [IJ:DSM:Sivasankari:2023, IJ:Scientometrics:Thierry:2023, IJ:KIS:STERGIOPOULOS:2024, IJ:AI:Thierry:2023]
@@ 3.2.3 AMiner
[IJ:ESWA:XIAO:2023] ? ur Papers: 3,499,032 Authors: 325,341
@@
[IJ:AS:Chen:2023] ? Papers: 500,000 Keywords: 9,668
[IJ:JOI:Yadav:2022] ASN + Scopus  [^13] https://github.com/Pratyush1296/SPACE-R Papers: 12,611 Authors: 22,603
@@
[IJ:mathematics:GUO:2024] ASN + AFT [^14] https://academictree.org/ Papers: 89,013 Students: 10,929
@@ 3.2.4 CiteULike
Among the works analyzed, 26.98% utilize some CiteULike version, yet these represent only 8.3% of the total datasets used, showing a greater reusability of the CiteULike dataset compared to others. In fact, almost all works that use CiteULike rely on CiteULike-a [J:EIJ:Huang:2024, IS:CSET:WU:2022, IJ:SC:XI:2024, IJ:CDS:AHMEDI:2022, IJ:ESWA:XI:2023, IC:MLBDBI:GUAN:2022, IJ:JS:Huang:2023, arXiv:KARIMI:2022, ARXIV:Mohamed:2022, IJ:JOI:Yadav:2022, IC:ICITACEE:Switrayana:2022] and CiteULike-t [J:EIJ:Huang:2024, IS:CSET:WU:2022, IJ:SC:XI:2024, IJ:ESWA:XI:2023, IC:MLBDBI:GUAN:2022, IJ:JS:Huang:2023, arXiv:KARIMI:2022, IC:ICITACEE:Switrayana:2022] without applying any filtering or modification.
@@
[IJ:ASC:Wang:2024, IJ:SMC:WANG:2022] ? ur Papers: 82,376 Users: 1,659
@@
[IC:KSEM:YU:2022] CiteULike-a Papers:11,845 Users: 4,880
@@
[IJ:IS:Chaudhuri:2022] CiteULike-a Papers: 1,000 Users: 10,327
@@
[IC:IKT:JAFARI:2021, IJ:AUT:Jafari:2023] CiteULike-a 171717ut-kdd/CiteUlike-SPIN-dataset(github.com) Papers: 16,980 Users: 5,551
@@ 3.2.5 Other source datasets
In addition to the previously mentioned datasets, there exists another group of publicly available datasets that, while published, are less frequently utilized in the analyzed works. Of the works analyzed, 17.46% utilize some of these datasets, representing 13.89% of the total datasets used; however, only three of these works employ the datasets in their original, unmodified form [IJ:SECS:Hamisu:2024, AC:ACIIDS:PAN:2024, IJ:EIT:Smail:2023].
@@
[IJ:SECS:Hamisu:2024, AC:ACIIDS:PAN:2024] Sugiyama [^18] http://scholarbank.nus.edu.sg/handle/10635/146027[DT:Sugiyama:2013, JCDL:Sugiyama:2013] Researchers: 50 Candidate Papers: 100,351
@@
[IC:ICDE:XIE:2022] US patent [^21] https://developer.uspto.gov/data Papers: 182,260 Authors: 73,974
@@
[IJ:IPM:Long:2024] PWC [^22] https://github.com/malteos/aspect-document-embeddings[IJ:JCDL:Ostendorff:2022] ur Papers: 88,940 Disciplines: Machine
@@
[IJ:IPM:Long:2024] SCICITE [^23] https://github.com/allenai/scicite ur Papers: 8,045 Disciplines:
@@
[IJ:ESWA:Wang:2023] Spotify [^24] https://www.kaggle.com/datasets/andrewmvd/spotify-playlists Papers: 11,703 Users: 1,720
[IJ:ESWA:Wang:2023] Yelp [^25] https://www.kaggle.com/datasets/yelp-dataset/yelp-dataset Papers: 19,479 Users: 10,918
[IJ:AS:Chen:2023] Metallurgical ? Papers: 5,702 Authors: 3,268
@@
[IJ:ESWA:MEI:2022] Arxiv HEP-TH [^26] https://snap.stanford.edu/data/cit-HepTh.html Papers: 18,066 Authors: 11,879
@@
[IJ:IEEE:SARWAR:2021] ? ur Disciplines: Computer Science
[IJ:IEEE:STERLING:2022] ? Papers: 2,271 Disciplines: Carbon,
@@
[IJ:Heliyon:LI:2024] DBLP Papers: 700
[IJ:AI:Xiao:2023] CNKI [^33] https://eval.cnki.net/index/ ur Disciplines:
@@
[IJ:AI:Xiao:2023] CNKI ur Disciplines: Data Mining
@@
[IJ:JOI:Yadav:2022] WoS  [^34] https://github.com/Pratyush1296/SPACE-R Papers: 27,442 Authors: 40,408
@@
[IJ:DATABASE:Kart:2022] PubMed  [^35] https://github.com/bioinfcollab/emati Disciplines:
@@
[IJ:ELECTRONICS:Blazevic:2023] Springer Nature
@@
[IJ:AMEC:KUS:2023] ArXiv Papers: 41,000 Users: 24,707
@@
[IJ:Scientometrics:GUNDOGAN:2022] ArXiv Papers: 122,825 Disciplines: Computer science
[IJ:Scientometrics:TAN:2023] ArXiv  [^36] https://github.com/opendata-ai/tr. Papers: 21,000 Years: 1991-2016
[IC:ICDE:XIE:2022] Scopus Papers: 1,304,907 Authors: 482,602
@@
[IJ:IS:Chaudhuri:2022] Scopus Size: 2000 Authors: 430
@@
[EC:ECIR:Takahashi:2022] Semantic Scholar Papers: 805,063 Disciplines: Engineering,
@@
[IC:WIST:Roßner:2023] Semantic Scholar +  [^37] https://opendata.iisys.de/opendata/Datasets/publication_similarity.zip Papers: 15,359 Publication Pairs:
@@
[IJ:IS:PINEDO:2024] Semantic Scholar +  [^38] https://github.com/IratxePinedoehu/arzigo Papers: 1795 Users: 736
@@ 3 Medicine
[IC:SEAI:Wang:2024] ACM Papers: 3,025 Authors: 5,835
@@
[IJ:Scientometrics:TAN:2023] USPTO  [^39] https://github.com/opendata-ai/tr. Papers: 21,000 Years: 2002-2018
@@ 3.3.1 What are the relevance criteria used to evaluate the recommendations?
44.29% of the works use references as a relevance criterion. 38.71% of them consider the author as the user, so the papers referenced by authors in their publications are used as a basis for relevance [IC:ASSP:Shen:2024, IC:DASFAA:Wang:2022, IC:EITCE:NIU:2023, IC:ICDE:XIE:2022, IJ:AI:GAO:2023, IJ:AS:Chen:2023, IJ:ESWA:Wang:2023, IJ:ESWA:XIAO:2023, IJ:IEEE:REN:2022, IJ:mathematics:GUO:2024, IJ:Scientometrics:Jiang:2023, IJ:TIIS:GUO:2024]. 38.71% of the works using references as relevance criteria rely on references found in the input paper [ARXIV:Shen:2024, IJ:AI:Thierry:2023, IJ:ESWA:MEI:2022, IJ:Heliyon:LI:2024, IJ:IEEE:HADHIATMA:2023, IJ:IEEE:STERLING:2022, IJ:JOI:Yadav:2022, IJ:Scientometrics:ALI:2022, IJ:Scientometrics:GUNDOGAN:2022, IJ:Scientometrics:Huang:2024, IJ:Scientometrics:Thierry:2023, IJ:Scientometrics:TAN:2023]. 12.9% of the works using references, even if the authors are not established as users, use datasets that only consider citation information for evaluating the system, so authors are considered the users, and their references are treated as interactions [PP:Mahesh:2023, IC:MLBDBI:GUAN:2022, IJ:IS:Chaudhuri:2022, IJ:ASC:Wang:2024]. 6.45% of the works use references from the same paper that is used to extract the input text [EC:ECIR:Takahashi:2022, IJ:DSM:Sivasankari:2023]. [IJ:AS:Wang:2024] not only considers positive examples the articles referenced by each author, but also the papers they have published.
@@
Out of the total works that use actual user interactions as a relevance criterion, 85% utilize the CiteULike dataset. The remaining works either use non-RPRS-specific datasets, such as Spotify or Yelp [IJ:ESWA:Wang:2023], or rely on artificially generated interactions [IJ:KIS:STERGIOPOULOS:2024, IJ:DATABASE:Kart:2022].
@@
12.86% of the works use some kind of label to identify relevant papers. These labels can be discipline classification labels that help match academic fields between users and recommended papers [CF:CSCW:LI:2022, IC:SEAI:Wang:2024] or between POI and recommended papers [IJ:AMEC:KUS:2023, IJ:EIT:Smail:2023]. Citation intent, which reflects the task or focus of the paper, is also used [IJ:IPM:Long:2024]. In addition, more explicit labels are sometimes applied to indicate relevance to particular authors [AC:ACIIDS:PAN:2024, AC:ACIIDS:PAN:2024] or topics [IJ:SHTI:GUO:2022].
@@
Only 8.57% of the works use human judgment to evaluate the recommendations obtained. This relevance criterion can help assess the recommendations by better aligning them with actual needs, leading to more accurate evaluations. The choice of criteria depends on the level of detail required and the specific context in which the evaluation is conducted. Some systems use simple binary ratings, such as categorizing publications as either relevant or irrelevant [J:JIS:MUNGEN:2022], or correct or incorrect with reasoning [IJ:IS:PINEDO:2024]. Others implement more nuanced scales, assigning points based on similarity or relevance strength [IC:WIST:Roßner:2023, IJ:ELECTRONICS:Blazevic:2023, IJ:IEEE:SARWAR:2021]. In addition to rating systems, some evaluations involve assessing the satisfaction of the users, highlighting the subjective nature of relevance judgment [IJ:Scientometrics:Kanwal:2024].
@@
Finally, alternative relevancy criteria in evaluating recommendations include assessing co-authorship potential by evaluating research compatibility [IJ:KBS:Xiao:2023], determining success based on the inclusion of papers by scientific research users [IJ:JES:LI:2024], and matching recommendations with user input parameters such as research topics and academic affiliations [IJ:AI:Xiao:2023].
@@
Paper references [ARXIV:Shen:2024, EC:ECIR:Takahashi:2022, IC:ASSP:Shen:2024, IC:DASFAA:Wang:2022, IC:EITCE:NIU:2023, IC:ICDE:XIE:2022, IC:MLBDBI:GUAN:2022, IJ:AI:GAO:2023, IJ:AI:Thierry:2023, IJ:AS:Chen:2023, IJ:AS:Wang:2024, IJ:ASC:Wang:2024, IJ:DSM:Sivasankari:2023, IJ:ESWA:MEI:2022, IJ:ESWA:Wang:2023, IJ:ESWA:XIAO:2023, IJ:Heliyon:LI:2024, IJ:IEEE:HADHIATMA:2023, IJ:IEEE:REN:2022, IJ:IEEE:STERLING:2022, IJ:IS:Chaudhuri:2022, IJ:JOI:Yadav:2022, IJ:mathematics:GUO:2024, IJ:Scientometrics:ALI:2022, IJ:Scientometrics:GUNDOGAN:2022, IJ:Scientometrics:Huang:2024, IJ:Scientometrics:Jiang:2023, IJ:Scientometrics:Thierry:2023, IJ:TIIS:GUO:2024, PP:Mahesh:2023, IJ:Scientometrics:TAN:2023] 44.29%
Interactions [arXiv:KARIMI:2022, ARXIV:Mohamed:2022, IC:ICITACEE:Switrayana:2022, IC:IKT:JAFARI:2021, IC:KSEM:YU:2022, IC:MLBDBI:GUAN:2022, IJ:ASC:Wang:2024, IJ:AUT:Jafari:2023, IJ:CDS:AHMEDI:2022, IJ:DATABASE:Kart:2022, IJ:ESWA:Wang:2023, IJ:ESWA:XI:2023, IJ:IS:Chaudhuri:2022, IJ:JOI:Yadav:2022, IJ:JS:Huang:2023, IJ:KIS:STERGIOPOULOS:2024, IJ:SC:XI:2024, IJ:SMC:WANG:2022, IS:CSET:WU:2022, J:EIJ:Huang:2024] 28.57%
Labeled Paper [AC:ACIIDS:PAN:2024, CF:CSCW:LI:2022, IC:SEAI:Wang:2024, IJ:AMEC:KUS:2023, IJ:EIT:Smail:2023, IJ:IPM:Long:2024, IJ:SECS:Hamisu:2024, IJ:SHTI:GUO:2022, IJ:Scientometrics:TAN:2023] 12.86%
Human judment [IC:WIST:Roßner:2023, IJ:ELECTRONICS:Blazevic:2023, IJ:IEEE:SARWAR:2021, IJ:IS:PINEDO:2024, IJ:Scientometrics:Kanwal:2024, J:JIS:MUNGEN:2022] 8.57%
Others [IJ:JES:LI:2024, IJ:KBS:Xiao:2023, IJ:AI:Xiao:2023] 4.29%